OpenStudy (anonymous):

Can someone please do this problem. Find k so that the line through (k-1,k+2) and (4,-1) is perpendicular to the line through (-3,2) and (2,5)

OpenStudy (anonymous):

Can anyone solve this problem it is part of my math project

OpenStudy (anonymous):

First find the gradient of the known line, which is (y2-y1)/(x2-x1) which comes out to 3/5. So we now know the gradient of the line we want is -5/3 because in order for the gradients to be perpendicular they must times together to = -1. We know now a gradient and a point on the line so we can work out its equation \[y-y1 = m(x-x1)\] which comes out to y+1 = -5/3(x-4) which --> 3y+3 = -5(x-4) --> 3y = -5x + 20 -3 --> y+5x = 17 k+2 + 5(k-1) = 17 k+2 + 5k -5 = 17 6k = 17 + 3 6k = 20 k = 6/20 k = 3/10 (Think I went wrong somewhere but can't work out where, might be right but seems an odd number).

OpenStudy (anonymous):

Where did you get 3y+3=-5(x-4) from

OpenStudy (anonymous):

I multipled through out by 3 to get rid of the fraction?

OpenStudy (anonymous):

Thanks this project is worth a test grade

OpenStudy (anonymous):

Oh dear, I really hope that is all correct for ya and GL.